MODS1 AGw Commissioning at LDG: 14 November 2010 UTC

Observer(s): R. Pogge (OSU), M. Pedani (LBTO)
Telescope Operator: S. Allanson (LBTO)
Telescope Support: M. Midkiff (LBTO)
Instrument Support: T. Shih (LBTO)
TCS testing team in Tucson: M. Delapena, N. Cushing, C. Biddick, et al.

Summary

Tonight was TCS testing of the BP6 binocular pointing build. At around midnight the TCS testing team asked us to run preset and guide/WFS tests with MODS1 at LDG. We continued in this activity from roughly 0700 to 1050UTC, at which time clouds rolled in and stopped observing.

After some bug fixes by the team in Tucson, we were able to command monocular presets with MODS in all relevant modes (track, acquire, guide, and active). Active optics collimation with the AGw WFS appeared to be working. Doug Miller came online late in the process to see how this was going. Seeing was variable, 0.7-1.5 arcsec, and we did not make any attempts to tune up any of the guide or WFS parameters since this was not the goal of the TCS testing.

MODS team verified that we are able to read out the data dictionary without problems for the telescope information of interest, so we do not have to make any immediate changes to our IIF agent code for MODS1.

At around 1000 UTC the TCS team finished their work and handed over the telescope to us for the rest of the night. We got about 45 minutes before heavy high clouds rolled in and stopped observing (if you can't see 8th mag on an 8.4m telescope, time to quit). Before getting shutdown we were able to re-measure the center-of-rotation of the MODS1 AGw guide camera after all the work on the ISS locking devices the previous week.

Details

BP6 Preset Tests with MODS1

0700 - The TCS testing crew in Tucson requested that we try presets with MODS1 using the BP6 TCS build.

Steven reconfigured the telescope at Zenith for MODS1 LDG operation (retracted the left tertiary swing arm). We updated the pointing and acquired a target (NGC1055, a galaxy nearby on the merdian and celestial equator).

Numerous bugs emerged with the guide acquire system going off. Michelle worked on this and found a solution that was coded in.

After this we were able to make guide presets.

We then tried active presets and were able to get reasonable WFS convergence. No runaways or other problems, but one of the first stars was well off axis and there was vignetting of the WFS image on one side (we may not be optimally centered in the WFS pickoff aperture - that is to be refined as part of tomorrow-night's official MODS1 AGw commissioning activities).

We were able to successfully perform active presets off-axis and on-axis, run various offsets (x y and rotator), and otherwise verify that we had achieved basic function needed to proceed with AGw commissioning observations.

We will confer during the day tomorrow among ourselves as to whether we want to try to run with BP6 as the "science" TCS build, or fall back on the AO5 build.

MODS1 AGw Guide Camera Center-of-Rotation

1000 - The TCS testing team in Tucson finished their work and handed over the telescope to us to finish out the night.

Setup on Persson star BS9118 (elevation about 56-deg). Did an active preset to tweak up the collimation for good image quality. Seeing was swimmy, about 1-1.5 arcsec.

1017 - Setup for a measurement of the Guide Camera center of rotation

Did a guide preset. Stopped guiding, drove the AGw guide probe to field center running open loop.

AGw configuration was as follows:

  • Guide Probe at nominal science-field center (GPROBE 0 0)

  • ND1.0 filter to cut down the bright light from the Persson star

  • Setup directory /OldRepository/20101114_MODS/ to hold guide camera snapshots taken with readGuideCamera

The LDG rotator was slewed to RotPos=0, took a reference image of 1s (left_guider000001.fits).

1020 - Slewed rotator 0..360-deg, and exposed guide cam 70sec.

Image: left_guider000002.fits

Exposure time of 70sec was too short, only got about half the arc. Signal level is good

1025 - Run 2: Slewed rotator 360..0-deg, exposed guide cam 120sec

Image: left_guider000003.fits

Got a complete ring, good signal.

The rotator gave a fault after the end of the slew. Steven stopped and restarted without problems.

1040 - snapped image for WFS hotspot position check

Issued a guide preset to put the Persson star back into the hotspot and start guiding. Then stopped guiding to take a snapshot with the star at the hotspot.

Image: left_guider000004.fits

Star is very faint - clouds are coming in.

1044 - Switching to a bright pointing star nearby: ACT0197 (7.1mag) - executing a guide preset.

Image: left_guider000005.fits

Good signal, but transparency and seeing starting to degrade. Before we could take a second image the star was gone. Outside we saw heavy overcast.

1050 - Closed, now completely overcast and cannot acquire stars.

Measurement of the C-o-R

Using image left_guider000003.fits, we measured the center of rotation using the ring bisector method:
   Xcor = 267.4
   Ycor = 212.3

With the GCS now reading out the acquisition images with the region of interest (subframe) of
  43 16 512 455
The guiding/WFS hot-spot is confirmed to be at
   Xwfs = 220
   Ywfs = 416
This will be checked on the next clear night to verify that we are optimally centered in the WFS aperture.

The GCS configuration file for MODS1/LDG currently has the following as the guide camera center for pointing:
   Xcen = 270
   Ycen = 225
This is slightly off from our measured center-of-rotation, by about 3 and 13 pixels (0.33 and 1.4 arcsec) in X and Y, respectively. This is less than 1mm in the focal plane proper.

We should assess how close is close enough for our purposes and decide how to proceed next.

  • MODS AGw Guide Camera, with reference positions plotted:
    gcam_ref.png
This image is a composite of the center-of-rotation image (left_guider000003.fits) and the WFS hotspot snapshot image (left_guider000005.fits). The regions plotted are:
   Cross - measured Center of Rotation
   Circled Cross - current guidecam_center_x/y coordinates in OSU_L.cfg
   Small Circle - guider hotspot position (measured and from OSU_L.cfg)
   Large Circle - approximate size of the 3-arcsec WFS aperture
   Square - outline of the 5x5-mm (8x8-arcsec) WFS pickoff beamsplitter cube
The beam splitter was mounted using epoxy which pulled slightly to one side on curing, which is why the beam splitter is not centered on the WFS aperture. Removing the beam splitter would likely destroy it so we have elected to adopt this off-center orientation.

-- RichardPogge - 14 Nov 2010
Topic attachments
I Attachment Action Size Date Who Comment
gcam_ref.pngpng gcam_ref.png manage 227 K 14 Nov 2010 - 22:18 RichardPogge MODS AGw Guide Camera, with reference positions plotted
Topic revision: r2 - 16 Aug 2011, RichardPogge
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback